# Breakdown of symmetry constraint in Floquet topological superconductor

### v1

@@ III Breakdown of symmetry constraint in Floquet system
where $\mathcal{B}_{pq}=\langle p,\textbf{k}|\nabla_{\textbf{k}}|q,\textbf{k}\rangle$ and $|p/q,\textbf{k}\rangle$ are the real eigenstates of $\mathcal{H}_{\text{eff}}({\bf k})$ under the reality requirement $P\mathcal{T}|p/q,\textbf{k}\rangle=|p/q,\textbf{k}\rangle$. It is straightforward to calculate the real Chern number $\nu=1$ and the multipole chiral number $N_{0}=0$ and $N_{\pi/T}=1$ for the parameter values in Fig. 2(a). Both the numerical quasienergy spectrum and the analytical topological invariants prove that a hybrid-order topological superconductor phases with the coexisting first-order Majorana boundary modes in the zero quasienergy gap and the second-order Majorana corner modes in the $\pi/T$ quasienergy gap are realized by the periodic driving. This result is similarly supported by the distributions of the zero- and $\pi/T$-quasienergy states in Figs. 2(c) and 2(d). Next, we further explore whether the first-order Majorana boundary modes can emerge in the $\pi/T$ quasienergy gaps [65]. The quasienergy spectra under the open boundary conditions in the two directions [see Figs. 2(e)] and only $y$ direction [see Figs. 2(f)] indicate the emergence of first-order gapless Majorana boundary modes in the $\pi/T$ quasienergy gap and second-order gapped Majorana corner modes in the zero quasienergy gap. They are well described by the calculated topological invariants $\nu=1$, $N_{0}=-2$, and $N_{\pi/T}=0$. The distribution of zero-quasienergy states in Fig. 2(g) confirms their second-order nature and the one of the $\pi/T$-quasienergy states in Fig. 2(h) verifies their first-order nature. Consequently, our Floquet-engineering scheme fundamentally breaks the symmetry constraints without altering the symmetries. Setting the system free from the symmetry constraint, the first-order topological superconductor phase is created in the CII class. This is impossible in static systems.
@@
We now investigate whether the two orders of topological superconductors can coexist in one quasienergy gap or not. It is remarkable to find from the quasienergy spectrum under the open boundary condition in Fig. 3(a) that the $\pi/T$ gap hosts two second-order gapped Majorana corner modes and the zero gap hosts the hybrid-order topological phases with coexisting first-order gapless Majorana boundary modes and second-order gapped Majorana corner modes. The calculation indicates that $N_{0}=-1$, $N_{\pi/T}=1$ and $\nu=1$, which explains the emergence of hybrid-order Majorana modes in the zero quasienergy gap and second-order one in the $\pi/T$ gap. The probability distributions in Figs. 3(c) and 3(d) further support this result. The quasienergy spectra in Figs. 3(e) and 3(f) reveal that the hybrid-order Majorana modes can also emerge in the $\pi/T$ gap. This result is well characterized by $N_{0}=2$, $N_{\pi/T}=1$, and $\nu=1$ and by the distributions in Figs. 3(g) and 3(h). Therefore, it can be concluded that within our periodic system, the first-order Majorana boundary modes can either exist independently or coexist with the second-order Majorana corner modes to form the hybrid-order topological superconductor.
@@ IV Discussion and conclusion
It is important to note that the step-like driving scheme used in our work is intended to guarantee that the periodic system maintains the symmetry classes as the static system. The driving scheme may be generalized to other schemes that satisfy this symmetry requirement. In recent years, simulation of Majorana modes has been achieved in various platforms, such as condensed matter systems [31, 32, 33, 34, 69, 27, 28] and cold atoms system [70, 71]. Rich topological phases induced by Floquet engineering have been realized in solid-state systems [72, 41], photonic system [73, 74, 75, 76, 77, 78], cold-atom system [79, 80], acoustic system [81], and Rydberg arrays system [82]. Simulating Floquet Majorana modes in superconducting qubits [58, 62] and cold-atom system [83] directly provides some feasible schemes for realizing Floquet Majorana modes in our work. These progresses provide strong experimental support to the realization of our result.
In summary, we have proposed a Floquet-engineering scheme to break the symmetry constraint of topological superconductors. It is found that, by applying a periodic driving on a topological superconductor system belonging to the symmetry classes BDI and CII, we can create first-order Majorana boundary modes without changing any symmetry. We have further discovered that, after releasing the system from the symmetry constraint, diverse exotic hybrid-order topological superconductor phases with coexisting first-order Majorana boundary modes and second-order Majorana corner modes not only in two different quasienergy gaps but also in one single gap are present in our periodically driven system. Refreshing one’s general belief on the classification rule of the topological superconductors set by symmetries, our Floquet-engineering scheme opens a door to explore exotic topological superconductors and prompts the simulation and application of Majorana modes, in particular with regard to overcoming the limitations set by unchangeable symmetries of the fabricated materials.
@@ V Acknowledgments
The work is supported by the National Natural Science Foundation of China (Grants No. 124B2090, No. 12275109, and No. 12247101), the Innovation Program for Quantum Science and Technology of China (Grant No. 2023ZD0300904), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2025-jdzx07), and the Natural Science Foundation of Gansu Province (No. 22JR5RA389 and No. 25JRRA799).

### v2

@@ III Breakdown of symmetry constraint in a Floquet system
where $(\nabla_{\textbf{k}}\times\mathcal{B})_{z}$ is the Berry curvature, and $\mathcal{B}_{pq}=\langle p,\textbf{k}|\nabla_{\textbf{k}}|q,\textbf{k}\rangle$ and $|p/q,\textbf{k}\rangle$ are the real eigenstates of $\mathcal{H}_{\text{eff}}({\bf k})$ under the reality requirement $P\mathcal{T}|p/q,\textbf{k}\rangle=|p/q,\textbf{k}\rangle$. Note that the conventional Chern number cannot describe the first-order topological phase in the $P\mathcal{T}$-symmetric system due to the boundary modes with opposite transport properties. Here, the real Chern number succeeds in characterizing this first-order topological phase by introducing the operator $\sigma_{y}\tau_{0}$ to distinguish the boundary modes with opposite transport properties. Therefore, the multipole chiral number and real Chern number utilize different symmetries and different bulk-band features obtained under distinct (real and momentum) spaces to characterize the second- and first-order topological phases, respectively. As long as the second- and first-order topological phases emerge in our system, they can be separately described by these independent topological invariants in a sufficient manner [77]. It is straightforward to calculate the real Chern number $\nu=1$ and the multipole chiral number $N_{0}=0$ and $N_{\pi/T}=1$ for the parameter values in Fig. 2(a). Because of the time-reversal symmetry $\mathcal{T}$ of our periodic system, the first-order boundary modes of our periodic system are helical boundary modes. Both the numerical quasienergy spectrum and the analytical topological invariants prove that a hybrid-order topological insulator phase with coexisting first-order helical boundary modes in the zero quasienergy gap and the second-order corner modes in the $\pi/T$ quasienergy gap are realized by periodic driving. This result is supported by the distributions of the zero- and $\pi/T$-quasienergy states in Figs. 2(c) and 2(d). Next, we further explore whether the first-order boundary modes can emerge in the $\pi/T$ quasienergy gaps [71]. The quasienergy spectra under the open boundary conditions in the two directions [see Figs. 2(e)] and only the $y$ direction [see Figs. 2(f)] indicate the emergence of first-order gapless helical boundary modes in the $\pi/T$ quasienergy gap and second-order gapped corner modes in the zero quasienergy gap. They are well described by the calculated topological invariants $\nu=1$, $N_{0}=-2$, and $N_{\pi/T}=0$. The distribution of zero-quasienergy states in Fig. 2(g) confirms their second-order nature and the one of the $\pi/T$-quasienergy states in Fig. 2(h) verifies their first-order nature. Since our static system in Fig. 2 does not exhibit either a first-order topological phase or a hybrid-order one, these exotic topological phases in our periodic system are all engineered by periodic driving. Consequently, our Floquet-engineering scheme fundamentally breaks the symmetry constraints without altering the symmetries. Setting the system free from the symmetry constraint, the first-order topological insulator phase is created in the CII class. This is impossible in static systems.
@@
We now investigate whether or not the two orders of topological insulators can coexist in one quasienergy gap. It is remarkable to find from the quasienergy spectrum under the open boundary condition in Fig. 3(a) that the $\pi/T$ gap hosts two second-order gapped corner modes and the zero gap hosts the hybrid-order topological phases with coexisting first-order gapless helical boundary modes and second-order gapped corner modes. The calculation indicates that $N_{0}=-1$, $N_{\pi/T}=1$ and $\nu=1$, which explains the emergence of hybrid-order topological modes in the zero quasienergy gap and the second-order one in the $\pi/T$ gap. The probability distributions in Figs. 3(c) and 3(d) further support this result. The quasienergy spectra in Figs. 3(e) and 3(f) reveal that the hybrid-order topological modes can also emerge in the $\pi/T$ gap. This result is well characterized by $N_{0}=2$, $N_{\pi/T}=1$, and $\nu=1$ and by the distributions in Figs. 3(g) and 3(h). Similar to the analysis in the preceding text, these exotic topological phases are all created by the periodic driving. It can be concluded that within our periodic system, the first-order helical boundary modes can either exist independently or coexist with the second-order corner modes to form a hybrid-order topological insulator.
@@ IV Discussion and conclusion
It is important to note that the step-like driving scheme used in our work is intended to guarantee that the periodic system maintains the symmetry classes as the static system. The driving scheme may be generalized to other schemes that satisfy this symmetry requirement. It is emphasized that the real Chern number Eq. (4) is defined by all occupied bands. It simultaneously contains information about both the zero and $\pi/T$ gaps. Therefore, we can only characterize the case where the first-order boundary modes exist either in the zero gap or in the $\pi/T$ gap. If first-order boundary modes simultaneously appear in both the zero and $\pi/T$ gaps, we currently have no means to characterize them. Here, we primarily reveal that periodic driving can overcome the symmetry constraints on topological phases. Thus, this insufficiency does not affect our main conclusion. In recent years, simulations of a topological insulator have been achieved in various platforms, such as condensed matter systems [23, 31, 34, 36], photonic systems [24, 25, 26, 29, 33], superconducting quantum processors [78], and Rydberg atom systems [79, 80, 81]. Rich topological insulator phases induced by Floquet engineering have been realized in solid-state systems [82, 43], photonic systems [83, 84, 85, 86, 87], cold-atom systems [88, 89] and acoustic systems [90]. This progress provides strong experimental support to the realization of our result.
In summary, we have proposed a Floquet-engineering scheme to break the symmetry constraint of topological insulators. It is found that, by applying a periodic driving on a topological insulator system belonging to the symmetry classes BDI and CII, we can create first-order helical boundary modes without changing any symmetry. We have further discovered that, after releasing the system from the symmetry constraint, diverse exotic hybrid-order topological insulator phases with coexisting first-order helical boundary modes and second-order corner modes not only in two different quasienergy gaps but also in one single gap are present in our periodically driven system. Refreshing one’s general belief on the classification rule of the topological insulators set by symmetries, our Floquet-engineering scheme opens a door to explore exotic topological insulators and prompts the simulation and application of topological modes, in particular with regard to overcoming the limitations set by unchangeable symmetries of the fabricated materials.
@@ V Acknowledgments
The work is supported by the National Natural Science Foundation of China (Grants No. 124B2090, No. 12275109, No. 92576202, and No. 12247101), the Innovation Program for Quantum Science and Technology of China (Grant No. 2023ZD0300904), the Fundamental Research Funds for the Central Universities (Grant No. lzujbky-2025-jdzx07), and the Natural Science Foundation of Gansu Province (No. 22JR5RA389 and No. 25JRRA799).